A coyote chasing a rabbit is moving 8.00 m/s due east at one moment and 8.80 m/s due south 3.95 s later. Let the x axis point due east and the y axis point due north.

Find the x and y components of the coyote’s average acceleration during that time.

Find the magnitude of the coyote’s average acceleration during that time.

Find the direction of the coyote’s average acceleration during that time
Answer #1

v,= 8.00 ( m) Is bina velecih C-j CS) 2-23 mls o-Y 二(12.3 己-2

> the person above has the right procedure but it needs some CORRECTIONS:
1) a_x = x/t which is -8.00 (east) not -8.80, that is y, so switch the right values
2) the angle is right, but because it is in the third quadrant, you have to add the previous 180° from x=0, so you have 42.3+180 = 222.3°

marta candia Mon, Feb 21, 2022 10:35 AM